What is the area of a square whose perimeter is 65 feet?
Answer by stanbon(75887)

What is the area of a square whose perimeter is 65 feet?
-----
side = 65/4 = 16 1/4
---
Area = (16.25)^2 = 264.0625
==============================
Cheers,
Stan H.
==========